OpenStudy (anonymous):

R(r^1+r^2) = r^1r^2, for r^2

OpenStudy (anonymous):

hmm no one seems to like this one ready?

OpenStudy (anonymous):

Ready

OpenStudy (anonymous):

\[R(r_1+r_2) = r_1r_2\]i am assuming they are subscripts right? not exponents

OpenStudy (anonymous):

i am assuming that because no one writes \(r^1\) but you do see \(r_1\)

OpenStudy (anonymous):

Oh yes I see the r1 not r^1 I'm sorry

OpenStudy (anonymous):

in any case the first step is to multiply out on the left using the distributive law and get \[Rr_1+Rr_2=r_1r_2\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

then lets put everything with an \(r_2\) in it on the left, and the other term on the right, via \[Rr_2-r_1r_2=-Rr_1\] i did two steps at one, subtrcted \(r_1r_2\) from both sides and also subtracted \(Rr_1\) from both sides

OpenStudy (anonymous):

is that step ok?

OpenStudy (anonymous):

if not let me know and i can explain further we have two more steps to go

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

ok now we want to get \(r_2\) by itself,so we have to factor it out of the expression on the left hand side. by the distributive law \(Rr_2-r_1r_2=(R-r_1)r_2\) we we can write \[(R-r_1)r_2=-Rr_1\]

OpenStudy (anonymous):

and finally to get \(r_2\) by itself we divide both sides by \(R-r_1\) to get \[r_2=\frac{-Rr_1}{R-r_1}\]

OpenStudy (anonymous):

or, if you prefer looking at fewer minus signs you would write this as \[r_2=\frac{Rr_1}{r_1-R}\]

OpenStudy (anonymous):

hope all steps are clear, i don't think i skipped any

OpenStudy (anonymous):

Thank you you helped a lot :)